OpenStudy (anonymous):

Solve for b: (2b)/5 + b/10 - 4 = 6 + 3b – 15

OpenStudy (anonymous):

The LCM or LCD of 5 and 10 is 10 So, (4b + b - 40)/10 = 3b - 9 Multiply by 10 both the sides, 5b - 40 = 30b - 90 25b = 50 b = 2 is the answer...